7 MOLES OF TETRA ATOMIC NON LINEAR GAS A AT 10 ATM AND T K ARE MIXED WITH 6 MOLES OF ANOTHER GAS B AT T/3 K AND 5 ATM IN A CLOSED RIGID VESSEL WITHOUT ENERGY TRANSFER WITH SORROUNDINGS IF THE FINAL TEMPRATURE OF THE MIXTURE WAS 5T/6 K , THEN GAS B IS?
DIATOMIC OR TRIATOMIC ?

Grade:11

1 Answers

Siledar Divakar
7 years ago
Cv of gas A = 3/2 R + 3/2 R + 6R = 9R
Let Cv of gas B = Cv
Heat lost by gas A = 7 x 9R (T- 5T/6)
Heat gained by gas B = 6 Cv (5T/6 – T/3)
63 R(T/6) = 6 Cv (T/2)
Cv=7 R/2
therefore, gas is diatomic = 3R/2 + R + R = 7 R/2